OpenStudy (anonymous):

Find f(-11/5) if f(n) = 5n + 6. -17 5 -5

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

lets see

OpenStudy (anonymous):

....

OpenStudy (anonymous):

-5

OpenStudy (anonymous):

thank youuuuu

OpenStudy (anonymous):

\[\LARGE \color{blue}{f(n)=5n+6}\] \[\LARGE \color{blue}{f(-\frac{11}{5})=5.(-\frac{11}{5})+6 = -11 +6 = -5}\]